What is the vapor pressure of water at 100 degree
Answer #1

13

answer : 760 mmHg at 100°C. For water,

=================================================

16. Silver crystallizes in a cubic closest packed structure. The radius of Ag atom is 1.44 Å (144 pm). Calculate the density of solid silver.

Density =m/v

Diagonal =r+2r+r= 4r

cell is face-centered.

d2 + d2 = (4r)2

2d2 =16 r2

d = sqrt of 8 r2

= r sqrt of 8

where r = the radius

r= 144 pm

d= = 144 sqrt of 8   =407 pm

volume = d^3= (407 pm)^3 =6.74 x10 ^7 pm^3

cubic centimetres =6.74 x10 ^7 pm^3 x (1 x10^-10 cm /pm)^3

= 6.74 x10 ^7 cm^23

Density =m/v

Plugging values in above equations

= (4 atoms ) (107.9 g/mol)(1 mol/6.023 x10623 atoms) /6.74 x10 ^7 cm^23

=10 .6 g/cm^3------answer